# Entropy-Conservative numerical fluxes for compressible Euler equations with thermally perfect gas models

### v1

@@ 1 Introduction
To achieve accurate computational fluid dynamics simulations, research has been devoted to devising numerical discretizations capable of automatically reproducing important features of the continuous equations, such as enforcing physical symmetries or conserving primary or secondary invariants of the considered equations. These schemes, usually referred to as structure-preserving methods, have been shown to provide increased robustness and reliability [1], and have been the subject of many studies in recent years ([2, 3, 4, 5, 6, 7, 8])
@@
In the context of compressible flow simulations, the use of Kinetic Energy Preserving (KEP) methods has been adapted from the incompressible case. These methods are the most established and widely used among structure-preserving schemes [9, 7, 10]. Another important class of methods is that based on Entropy Conservative (EC) discretizations [11, 12, 13], which guarantee a correct induced discrete balance of entropy in absence of discontinuities. A fundamental difference between the two formulations is that, unlike KEP schemes, which only depend on the discretization of the continuity and momentum equations, EC schemes depend on the discretization of the energy equation, which, in turn, depends on the Equation of State (EoS) assumed for the gas.
@@
In the case of a calorically perfect gas, several methods have been developed that can preserve the entropy balance at the discrete level both exactly ([11, 14, 13, 15]) or approximately with arbitrary accuracy ([16, 17, 18]). In recent years, some extensions to the case of thermally perfect gases have also been developed [19, 20, 21]. The use of this alternative gas model more accurately describes the behavior at the high temperatures typical of combustion phenomena [22] as well as many wall-bounded high-speed flows, when temperature may exhibit both high absolute values and wide ranges [23, 24, 25].
@@
In a recent paper [26], the present authors developed a quite general methodology for the design of EC schemes for the system of compressible Euler equations with an arbitrary EoS, which reduces to already known methods in the case of a calorically perfect gas. Being completely general, the methodology can be easily applied to the case of a thermally perfect gas as well, which is a special case in which the perfect-gas EoS is still used in conjunction with an arbitrary relation between internal energy and absolute temperature only. In this note, we show how the peculiarities of the thermally perfect gas model can be exploited to particularize the general formulation, yielding EC discretizations suitable for smooth regions of the flow. Specifically, this approach alleviates the intrinsic potential singularity of the original method and obtains more robust and accurate formulations, as compared to existing ones.
@@ 2 Problem formulation
where $\mathbf{u}$ and $\mathbf{f}$ are the variables and flux vectors, respectively. For the single-component case one has $\mathbf{u}=(\rho,\rho u,\rho E)^{\text{T}}$ and $\mathbf{f}=(\rho u,\rho uu+p,u(\rho E+p))^{\text{T}}$ where $\rho$ and $\rho u$ are the density and momentum ($u$ being the velocity), $p$ is the pressure, and $E$ the total energy per unit mass, given by the sum of kinetic energy $\kappa=u^{2}/2$ and internal energy $e$: $E=\kappa+e$. The one-dimensional case will be used for ease of exposition, but the theory presented here can be straightforwardly extended to the multi-dimensional case. The thermally perfect gas model will be considered in this study, meaning that the usual perfect-gas EoS will be assumed: $p=\rho RT$, where $R$ is the gas constant and $T$ is the absolute temperature, but the isochoric specific heat capacity $c_{v}$ is dependent on temperature, which implies that internal energy can be expressed as $e=\int_{T_{\mathrm{ref}}}^{T}c_{v}(T^{\prime})\,\mathrm{d}{T^{\prime}}+e_{\text{ref}}$ and the “ref" subscript indicates some reference condition. Mayer’s relation $c_{p}-c_{v}=R$ is still valid, $c_{p}$ being the isobaric specific heat.
@@
In Eq. (2) the convective fluxes for mass and internal energy are still unspecified, and are residual degrees of freedom at our disposal to enforce additional structural properties to the formulation. The objective of the present treatment is to derive a KEP discretization which is also conservative of entropy, defined for the thermally perfect gas as $s=\int_{T_{\mathrm{ref}}}^{T}\frac{c_{v}(T^{\prime})}{T^{\prime}}\,\mathrm{d}{T^{\prime}}-R\log{(\rho/\rho_{\mathrm{ref}}})+s_{\mathrm{ref}}$, and with favourable properties in terms of robustness and accuracy.
@@ 2.1 Derivation of the fluxes
Eq. (3) makes use of the difference operator $\delta^{\,+}\phi=\phi_{i+1}-\phi_{i}$ and introduces the Gibbs free energy $g=h-Ts$, with $h=e+p/\rho$ the specific enthalpy. Eq. (3) still does not constrain the mass flux $\mathcal{F}_{\rho}$, leaving it free and giving rise to a family of EC discretizations, but it has the drawback of becoming singular in the case of constant temperature. For thermally perfect gases one has $g=e+T(R-s)$, leading to
$g=\omega(T)+RT\log{\rho};\quad\quad\quad\omega(T)=\int_{T_{\mathrm{ref}}}^{T}c_{v}(T^{\prime})\,\mathrm{d}{T^{\prime}}-\int_{T_{\mathrm{ref}}}^{T}\frac{c_{v}(T^{\prime})}{T^{\prime}}\,\mathrm{d}{T^{\prime}}+T(R-R\log\rho_{\mathrm{ref}}-s_{\mathrm{ref}})+e_{\mathrm{ref}},$ (4)
which separates the term with dependence on density from those dependent on temperature only. By substituting Eq. (4) in Eq. (3), it is possible to see that the dependence of the internal-energy numerical flux on pressure can be eliminated with a suitable choice of the mass flux. In particular, the choice $\mathcal{F}_{\rho}=\overline{\rho}^{\text{log}}\overline{u}\,$, where $\overline{\phi}^{\text{log}}=\delta^{\,+}\phi/\delta^{\,+}\log\phi$ is the logarithmic mean, nullifies the pressure term in Eq. (3) and the final form for the convective flux of internal energy guaranteeing entropy conservation becomes
@@
where the flux correctly depends only on temperature. If the calorically perfect gas model is considered in place of the thermally perfect one, Eq. (5) and (2) translate into the EC flux of Ranocha [13] (see Eq. (4) below). For the thermally perfect gas, progress can be made only by assuming a specific functional dependence $c_{v}(T)$.
@@
${c}_{v}(T)=\sum_{m=-L}^{R}c_{m}{T}^{m}.$ (6)
We will detail the derivation of the EC fluxes for this general formulation, even though usually only 5, 7 or 9 coefficients are used to experimentally fit the gas behavior [32]. In addition to being widely used, this form of $c_{v}(T)$ is also easily tractable from an analytical point of view. For another example of functional dependence $c_{v}(T)$, see Sec. 3.1.
@@
$\dfrac{\delta^{\,+}\omega/T}{\delta^{\,+}1/T}=e_{\mathrm{ref}}+c_{-1}\left(1+\dfrac{\delta^{\,+}\frac{\log{T/T_{\mathrm{ref}}}}{T}}{\delta^{\,+}1/T}\right)+c_{0}\dfrac{\delta^{\,+}\log{1/T}}{\delta^{\,+}1/T}-\sum_{\begin{subarray}{c}m=-L\\ m\neq-1,0\end{subarray}}^{R}\frac{c_{m}}{m(m+1)}\frac{\delta^{\,+}T^{m}}{\delta^{\,+}1/T}.$
The variable term associated with $c_{-1}$ can be manipulated as
$\dfrac{\delta^{\,+}\frac{\log{T/T_{\mathrm{ref}}}}{T}}{\delta^{\,+}1/T}=-\dfrac{\overline{1/T}\,}{\overline{1/T}^{\text{log}}}+\overline{\log T/T_{\mathrm{ref}}}\,,$
@@
$a^{k}-b^{k}=-\left(\frac{1}{a}-\frac{1}{b}\right)\sum_{\ell=1}^{k}a^{k-\ell+1}b^{\ell}=(a-b)\sum_{\ell=0}^{k-1}a^{k-\ell-1}b^{\ell},$ (7)
@@
$\dfrac{\delta^{\,+}T^{m}}{\delta^{\,+}1/T}=-\left(\sum_{\ell=1}^{m}T_{i+1}^{m-\ell+1}T_{i}^{\ell}\right)\qquad m\geq 1,$
@@
$\dfrac{\delta^{\,+}T^{m}}{\delta^{\,+}1/T}=\left(\sum_{\ell=m+1}^{0}T_{i+1}^{m-\ell+1}T_{i}^{\ell}\right)\qquad m\leq-2.$
@@
$\overline{\overline{T}}^{\,m}=\dfrac{\sum_{\ell=1}^{m}T_{i+1}^{m-\ell+1}T_{i}^{\ell}}{m}\quad\text{for }m\geq 1\quad\quad\quad\text{and}\quad\quad\quad\widetilde{\widetilde{T}}^{\,m}=\dfrac{\sum_{\ell=m+1}^{0}T_{i+1}^{m-\ell+1}T_{i}^{\ell}}{|m|}\quad\text{for }m\leq-2,$
[…]
$\mathcal{F}^{\text{c}}_{\rho e}=\mathcal{F}_{\rho}\left[e_{\mathrm{ref}}+c_{-1}\left(1-\dfrac{\overline{1/T}\,}{\overline{1/T}^{\text{log}}}+\overline{\log T/T_{\mathrm{ref}}}\,\right)+c_{0}\dfrac{1}{\overline{1/T}^{\text{log}}}+\sum_{\begin{subarray}{c}m=-L\end{subarray}}^{-2}\frac{c_{m}}{m+1}\widetilde{\widetilde{T}}^{m}+\sum_{\begin{subarray}{c}m=1\end{subarray}}^{R}\frac{c_{m}}{m+1}\overline{\overline{T}}^{m}\right].$ (8)
The scheme based on Eqs. (2), (5) and (8) is exactly entropy conserving. The use of the mass flux $\mathcal{F}_{\rho}=\overline{\rho}^{\text{log}}\overline{u}\,$ has reduced, in the case of thermally perfect gases, the singularity of the original formulation for an arbitrary EoS to that associated with the computation of the logarithmic mean, for which a local fix analogous to that devised by Ismail and Roe [12] can be used. Asymptotycally Entropy Conserving (AEC) formulations similar to that developed by De Michele and Coppola [16] could also be easily derived by using the asymptotic expansion of the logarithm employed in the calorically perfect case. In fact, using the Taylor series expansion
@@
$\mathcal{F}_{\rho e}^{\text{c}\,\text{AEC}(\!N\!)}=\mathcal{F}_{\rho}^{\text{AEC}(\!N\!)}\!\!\left[e_{\mathrm{ref}}+\left(c_{0}\overline{T}^{H}\!\!-c_{-1}\right)\sum_{n=0}^{N}\frac{\hat{T}^{2n}}{2n+1}+c_{-1}\left(1+\overline{\log{T/T_{\mathrm{ref}}}}\,\right)\!+\!\!\!\!\sum_{\begin{subarray}{c}m=-L\end{subarray}}^{-2}\frac{c_{m}}{m+1}\widetilde{\widetilde{T}}^{m}\!\!\!+\!\!\sum_{\begin{subarray}{c}m=1\end{subarray}}^{R}\frac{c_{m}}{m+1}\overline{\overline{T}}^{m}\right]\!.$ (10)
As in the case of calorically perfect gases, this set of fluxes constitutes a non-singular hierarchy of discretizations mainly based on simple algebraic operations and with increasing EC property, converging to Eq. (5) and (8) as $N\to\infty.$

### v2

# Formulation of entropy-conservative discretizations for compressible flows of thermally perfect gases

@@ 1 Introduction
Accurate and robust numerical simulations play an invaluable role in the study of turbulent compressible flows; however, standard discretizations of the equations are known to suffer from nonlinear instabilities. This is particularly pronounced at high Reynolds numbers, even in the absence of shock waves, and it is often attributed to the use of nondissipative schemes for the approximation of the convective terms [1].
@@
To address these challenges, research has been devoted to devising numerical discretizations capable of automatically reproducing important features of the continuous equations, such as enforcing physical symmetries or conserving primary or secondary invariants of the considered equations [2]. These schemes, usually referred to as structure-preserving methods, have been shown to provide increased robustness and reliability, and have been the subject of many studies in recent years ([3, 4, 5, 6, 7, 8, 9])
In incompressible flows, kinetic energy acts as a bounded $L_{2}$ measure of velocity [10], and as such it acts as a mathematical entropy for the incompressible Euler equations and can be used to assess nonlinear stability. In the context of compressible flow simulations, kinetic energy is no longer a conserved quantity and cannot provide a bound on the solution, but still the use of Kinetic Energy Preserving (KEP) methods has been successfully adapted from the incompressible case. These methods are the most established and widely used among structure-preserving schemes [11, 12, 8]. The conditions to obtain KEP schemes have been defined both in the context of finite volume methods [13, 14] and finite difference methods [8, 15]. However, the most commonly used definition for the KEP property introduces some indeterminacy, due to the arbitrariness of the discretization of the pressure term  [16, 17], which nevertheless has been found to have a significant impact on the evolution of kinetic energy [18].
@@
In compressible flows, kinetic energy no longer serves as a mathematical entropy of the governing equations; instead, in the inviscid limit, this role is primarily assumed by the thermodynamic entropy, whose existence is closely linked to the symmetrization of the Euler equations [19, 20]. Accordingly, another important class of methods is that based on Entropy Conservative (EC) discretizations [21, 22, 23], which guarantee a correct induced discrete balance of entropy in absence of discontinuities. A fundamental difference between the two formulations is that, unlike KEP schemes, which only depend on the discretization of the continuity and momentum equations, EC schemes depend on the discretization of the energy equation, which, in turn, depends on the Equation of State (EoS) assumed for the gas.
In the case of a calorically perfect gas, several methods have been developed that can preserve the entropy balance at the discrete level both exactly ([21, 24, 23, 25]) or approximately with arbitrary accuracy ([26, 27, 28]). Most EC schemes have been developed within the framework of the technique introduced by Tadmor [29], which has been applied to finite volume and finite element schemes. An improvement in the affordability of EC schemes came from the work of Ismail and Roe [22], which introduced the use of fluxes based on the logarithmic mean serving as the foundation of many subsequent developments. Building on this, Chandrashekar [24] proposed the first scheme that simultaneously satisfied both the KEP and the EC property. More recently, Ranocha [23, 30, 16] introduced a different KEP and EC formulation which, among other advantages, was shown to have a discretization of the pressure which led to more physically consistent treatment of kinetic energy [18].
@@
An alternative gas model is that of the thermally perfect gas, in which it is taken into account the temperature variation of specific heats caused by the excitation of vibrational energy within the molecules of the gas and by the electronic energy associated with electron motion within the atoms and molecules [31]. The use of this model more accurately describes the behavior at the high temperatures typical of combustion phenomena [32] as well as many wall-bounded high-speed flows, when temperature may exhibit both high absolute values and wide ranges [33, 34, 35]. In recent years, some extensions of the EC schemes to the case of thermally perfect gases have also been developed [36, 37, 38]. Notably, all these formulations reduce to the scheme of Chandrashekar [24] in the limit of a calorically perfect gas, owing to an analogous treatment of the pressure terms.
@@
In a recent paper [39], the present authors developed a quite general methodology for the design of KEP and EC schemes for the system of compressible Euler equations with an arbitrary EoS, which can be reduced to the already known method of Ranocha [23] in the case of a calorically perfect gas. Being completely general, the methodology can be easily applied to the case of a thermally perfect gas as well, which is a special case in which the perfect-gas EoS is still used in conjunction with an arbitrary relation between internal energy and absolute temperature only. However, in its original formulation, the scheme exhibited a singularity in the limit of a constant temperature field, necessitating the local use of an alternative, nonsingular flux in such cases.
In this paper, we show how the peculiarities of the thermally perfect gas model can be exploited to particularize the general formulation developed for an arbitrary EoS, yielding EC discretizations suitable for smooth regions of the flow. Specifically, this approach alleviates the intrinsic potential singularity of the original method. Moreover, due to the difference in the treatment of the pressure terms, the formulation obtained is more robust and physically consistent as compared to existing ones.
The structure of the paper is as follows. In Section 2, we introduce the governing equations, including the thermodynamic closure for thermally perfect gases. We then recall the necessary conditions for kinetic energy preservation and entropy conservation, and show how the general EC condition particularizes to this case. A polynomial-based model for the temperature dependence of specific heats is adopted, leading to a discrete formulation that is EC, KEP, and free of singularities. A comparison with existing schemes is also presented.
Section 3 extends the formulation to a different internal energy model, i.e. the Rigid-Rotor Harmonic-Oscillator (RRHO) model; we also generalize the scheme to multicomponent, nonreacting mixtures under a single temperature assumption.
Numerical results are provided in Section 4 to assess the conservation properties and robustness of the schemes on two benchmark problems: an inviscid doubly periodic jet, and the Taylor–Green vortex.
Finally, Section 5 presents concluding remarks, summarizes the findings, and outlines future research directions
@@ 2 Problem formulation
where $\mathbf{u}$ and $\mathbf{f}$ are the variables and flux vectors, respectively. For the single-component case one has $\mathbf{u}=(\rho,\rho u,\rho E)^{\text{T}}$ and $\mathbf{f}=(\rho u,\rho uu+p,u(\rho E+p))^{\text{T}}$ where $\rho$ and $\rho u$ are the density and momentum ($u$ being the velocity), $p$ is the pressure, and $E$ the total energy per unit mass, given by the sum of kinetic energy $\kappa=u^{2}/2$ and internal energy $e$: $E=\kappa+e$. The one-dimensional case will be used for ease of exposition, but the theory presented here can be straightforwardly extended to the multidimensional case. The thermally perfect gas model will be considered in this study, meaning that the usual perfect-gas EoS will be assumed: $p=\rho RT$, where $R$ is the gas constant and $T$ is the absolute temperature, but the isochoric specific heat capacity $c_{v}$ is dependent on temperature, which implies that internal energy can be expressed as
$e=\int_{T_{\mathrm{ref}}}^{T}c_{v}(T^{\prime})\,\mathrm{d}{T^{\prime}}+e_{\text{ref}}$
and the “ref" subscript indicates some reference condition. Mayer’s relation $c_{p}-c_{v}=R$ is still valid, $c_{p}$ being the isobaric specific heat.
@@
In Eq. (2) the convective fluxes for mass and internal energy are still unspecified, and are residual degrees of freedom at our disposal to enforce additional structural properties to the formulation. The objective of the present treatment is to derive a KEP discretization which is also conservative of entropy, defined for the thermally perfect gas as
$s=\int_{T_{\mathrm{ref}}}^{T}\frac{c_{v}(T^{\prime})}{T^{\prime}}\,\mathrm{d}{T^{\prime}}-R\log{(\rho/\rho_{\mathrm{ref}}})+s_{\mathrm{ref}},$
and with favourable properties in terms of robustness and accuracy.
@@ 2.1 Derivation of the fluxes
Eq. (3) makes use of the difference operator $\delta^{\,+}\phi=\phi_{i+1}-\phi_{i}$ and introduces the Gibbs free energy g=hTsg=h-Ts, with $h=e+p/\rho$ the specific enthalpy. Eq. (3) still does not constrain the mass flux $\mathcal{F}_{\rho}$, leaving it free and giving rise to a family of EC discretizations, but it has the drawback of becoming singular in the case of constant temperature. For thermally perfect gases one has g=e+T(Rs)g=e+T(R-s), leading to
$g=\omega(T)+RT\log{\rho};\quad\quad\quad\omega(T)=\int_{T_{\mathrm{ref}}}^{T}c_{v}(T^{\prime})\,\mathrm{d}{T^{\prime}}-T\int_{T_{\mathrm{ref}}}^{T}\frac{c_{v}(T^{\prime})}{T^{\prime}}\,\mathrm{d}{T^{\prime}}+T(R-R\log\rho_{\mathrm{ref}}-s_{\mathrm{ref}})+e_{\mathrm{ref}},$ (4)
which separates the term with dependence on density from those dependent on temperature only. By substituting Eq. (4) in Eq. (3) one has
$\mathcal{F}^{\text{c}}_{\rho e}=\mathcal{F}_{\rho}\,\dfrac{\delta^{\,+}\omega/T}{\delta^{\,+}1/T}+\mathcal{F}_{\rho}\,R\dfrac{\delta^{\,+}\log\rho}{\delta^{\,+}1/T}-\overline{u}\,\,\dfrac{\delta^{\,+}p/T}{\delta^{\,+}1/T}.$
which, by using the perfect gas EoS can be rearranged as
$\mathcal{F}^{\text{c}}_{\rho e}=\mathcal{F}_{\rho}\,\dfrac{\delta^{\,+}\omega/T}{\delta^{\,+}1/T}+\,R\dfrac{\delta^{\,+}\log\rho}{\delta^{\,+}1/T}\left(\mathcal{F}_{\rho}-\overline{u}\,\,\dfrac{\delta^{\,+}\rho}{\delta^{\,+}\log\rho}\right).$ (5)
From Eq. (5) it is possible to see that the dependence of the internal-energy numerical flux on pressure can be eliminated with a suitable choice of the mass flux. In particular, the choice $\mathcal{F}_{\rho}=\overline{\rho}^{\text{log}}\overline{u}\,$, where $\overline{\phi}^{\text{log}}=\delta^{\,+}\phi/\delta^{\,+}\log\phi$ is the logarithmic mean, nullifies the second term at the r.h.s. of Eq. (5) and the final form for the convective flux of internal energy guaranteeing entropy conservation becomes
@@
where now the internal-energy flux is built as the product between the mass flux and a term constituting a consistent average of internal energy, which correctly depends only on temperature. If the calorically perfect gas model is considered in place of the thermally perfect one, Eq. (6) and (2) translate into the EC flux of Ranocha [23] (see Eq. (21) below). For the thermally perfect gas, progress can be made only by assuming a specific functional dependence $c_{v}(T)$.
2.2 Polynomial-based fitting model
@@
${c}_{v}(T)=\sum_{m=-\ell}^{r}c_{m}{T}^{m}.$ (7)
In Eq. (7) $\ell$ and $r$ are arbitrary natural numbers, whose values are used to give flexibility to the fitting capability of the functional dependence $c_{v}(T)$. We will detail the derivation of the EC fluxes for this general formulation, even though usually only 5, 7 or 9 coefficients are used to experimentally fit the gas behavior [42, 43]. In addition to being widely used, this form of $c_{v}(T)$ is also easily tractable from an analytical point of view. For another example of functional dependence $c_{v}(T)$, see Sec. 3.1.
@@
$\dfrac{\delta^{\,+}\omega/T}{\delta^{\,+}1/T}=\varepsilon_{\mathrm{ref}}+c_{-1}\left(1+\dfrac{\delta^{\,+}\frac{\log{T}}{T}}{\delta^{\,+}1/T}\right)+c_{0}\dfrac{\delta^{\,+}\log{1/T}}{\delta^{\,+}1/T}-\sum_{\begin{subarray}{c}m=-\ell\\ m\neq-1,0\end{subarray}}^{r}\frac{c_{m}}{m(m+1)}\frac{\delta^{\,+}T^{m}}{\delta^{\,+}1/T}.$ (8)
in which
$\varepsilon_{\mathrm{ref}}=e_{\mathrm{ref}}-c_{-1}\log{T_{\mathrm{ref}}}-\sum_{\begin{subarray}{c}m=-\ell\\ m\neq-1\end{subarray}}^{r}\frac{c_{m}}{m+1}T_{\mathrm{ref}}^{m+1}$
and it is equivalent to the $b_{1}$ coefficient present in NASA polynomials in [43]. The variable term associated with $c_{-1}$ in Eq. (8) can be manipulated as
$\dfrac{\delta^{\,+}\frac{\log{T}}{T}}{\delta^{\,+}1/T}=-\dfrac{\overline{1/T}\,}{\overline{1/T}^{\text{log}}}+\overline{\log T}\,$,
@@
$a^{k}-b^{k}=-\left(\frac{1}{a}-\frac{1}{b}\right)\sum_{\nu=1}^{k}a^{k-\nu+1}b^{\nu}=(a-b)\sum_{\nu=0}^{k-1}a^{k-\nu-1}b^{\nu},$ (9)
@@
$\dfrac{\delta^{\,+}T^{m}}{\delta^{\,+}1/T}=-\left(\sum_{\nu=1}^{m}T_{i+1}^{m-\nu+1}T_{i}^{\nu}\right)\qquad m\geq 1,$
@@
$\dfrac{\delta^{\,+}T^{m}}{\delta^{\,+}1/T}=\left(\sum_{\nu=m+1}^{0}T_{i+1}^{m-\nu+1}T_{i}^{\nu}\right)\qquad m\leq-2.$
@@
$\overline{\overline{T}}^{\,m}=\dfrac{\sum_{\nu=1}^{m}T_{i+1}^{m-\nu+1}T_{i}^{\nu}}{m}\quad\text{for }m\geq 1\quad\quad\quad\text{and}\quad\quad\quad\widetilde{\widetilde{T}}^{\,m}=\dfrac{\sum_{\nu=m+1}^{0}T_{i+1}^{m-\nu+1}T_{i}^{\nu}}{|m|}\quad\text{for }m\leq-2,$
both of which are consistent with $T^{m+1}$, with $\overline{\overline{T}}^{\,1}=\overline{\overline{\left(T,T\right)}}$ and $\widetilde{\widetilde{T}}^{\,-2}=\overline{\left(1/T\right)}$. The final result is
$\mathcal{F}^{\text{c}}_{\rho e}=\mathcal{F}_{\rho}\left[e_{\mathrm{ref}}+c_{-1}\left(1-\dfrac{\overline{1/T}\,}{\overline{1/T}^{\text{log}}}+\overline{\log T}\,\right)+c_{0}\dfrac{1}{\overline{1/T}^{\text{log}}}+\sum_{\begin{subarray}{c}m=-\ell\end{subarray}}^{-2}\frac{c_{m}}{m+1}\widetilde{\widetilde{T}}^{m}+\sum_{\begin{subarray}{c}m=1\end{subarray}}^{r}\frac{c_{m}}{m+1}\overline{\overline{T}}^{m}\right].$ (10)
The scheme based on Eqs. (2), (6) and (10) is exactly entropy conserving. The use of the mass flux $\mathcal{F}_{\rho}=\overline{\rho}^{\text{log}}\overline{u}\,$ has reduced, in the case of thermally perfect gases, the singularity of the original formulation for an arbitrary EoS to that associated with the computation of the logarithmic mean, for which a local fix analogous to that devised by Ismail and Roe [22] can be used. Asymptotically Entropy Conserving (AEC) formulations similar to that developed by De Michele and Coppola [26] could also be easily derived by using the asymptotic expansion of the logarithm employed in the calorically perfect case. In fact, using the Taylor series expansion
@@
$\mathcal{F}_{\rho e}^{\text{c AEC}(N)}=\mathcal{F}_{\rho}^{\text{AEC}(N)}\left[\varepsilon_{\mathrm{ref}}+\left(c_{0}\overline{T}^{H}-c_{-1}\right)\sum_{n=0}^{N}\frac{\hat{T}^{2n}}{2n+1}+c_{-1}\left(1+\overline{\log{T}}\,\right)+\sum_{\begin{subarray}{c}m=-\ell\end{subarray}}^{-2}\frac{c_{m}}{m+1}\widetilde{\widetilde{T}}^{m}+\sum_{\begin{subarray}{c}m=1\end{subarray}}^{r}\frac{c_{m}}{m+1}\overline{\overline{T}}^{m}\right],$ (12)
where $\overline{\phi}^{H}$ is the harmonic mean $\overline{\phi}^{H}=\left(\overline{1/\phi}\,\right)^{-1}$. As in the case of calorically perfect gases, this set of fluxes constitutes a nonsingular hierarchy of discretizations mainly based on simple algebraic operations and with increasing EC property, converging to Eq. (6) and (10) as $N\to\infty.$